Fermi National Laboratory


Accelerator Update for November 7, 2005 - November 9, 2005

Monday November 7

The day shift began with Operations monitoring store 4490 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with Booster experts conducting studies, with MiniBooNE off due to an upcoming access, and with NuMI and the Meson MCenter experiments taking beam.

At 8 AM, the P1 transfer line suffered a vacuum burst.

Operations began preparing the Recycler to stash at 8:06 AM. The stash ended. MCenter resumed taking beam at 8:56 AM.

Pbar experts began target tuning at 9:24 AM.

A Main Injector (MI) expert began a study period at 11:30 AM.

Booster studies ended at 12:14 PM.

The MiniBooNE LCW work finished at 2:32 PM. Operations established beam about twenty minutes later.

At 4:09 PM, the Tevatron (TeV) suffered a sector F4 quench due to a TeV RF station (TRF4) sparking. Store 4490 was lost. An RF expert investigated.

The RF expert reported at 5:32 PM that they found nothing wrong with TRF4, but they did find and replace a bad attenuator on TRF1.

Operations established beam to the Meson MTest experiment at 6:31 PM.

At 6:42 PM, a TeV F sector kicker magnet (T:F17K3) alarmed on an overtemp indication.

At 7:22 PM, Electron Cooling experts reported that they were having a problem with the Pelletron and it would be down for a while.

Operations prepared the TeV F sector for access at 7:25 PM to investigate the kicker cooling problem.

E-Cool experts reported at 8:46 PM that the Pelletron would have to be drained and refilled with its cooling fluid, a six-hour process. They will begin this job on Tuesday morning.

The TeV kicker experts reported at 8:52 PM that they had filled the kicker reservoirs. The alarm cleared.

Operations reestablished beam to Pbar, MiniBooNE, and NuMI at 9:42 PM.

Operations reestablished beam to MCenter at 9:54 PM.

Operations put the TeV through a dry squeeze at 10:27 PM.

Operations put the TeV into shot setup at 10:56 PM.

Tuesday November 8

The midnight shift began with Operations monitoring the stash, with the TeV in shot setup, with Pbar and MCenter off due to the shot setup, and with MiniBooNE and NuMI taking beam.

Operations established Recycler only store 4492 at 1:22 AM with an initial luminosity of 120.6E30. Pbar, MTest, and MCenter resumed normal operations.

Operations began preparing the Recycler to stash at 2:06 AM. The stash ended. Pbar, MTest, and MCenter resumed normal operations at 3:16 AM.

Recycler began preparing to stash at 6:44 AM. The stash ended. Pbar, MTest, and MCenter resumed normal operations at 7:37 AM.

Operations reported at 8:44 AM that TeV experts were beginning the TeV sector A1 BPM upgrade. The upgrade took about four hours to complete.

At 9:32 AM, Operations reported that NuMI was tripping off due to bad beam position. Operators informed experts and turned off their beam.

A MiniBooNE expert turned their Horn’s current down to half its normal level for a test at 9:51 AM.

Proton experts began a reverse proton study at 10:40 AM.

Operations resumed sending beam to NuMI at 10:53 AM, but at a lower intensity. They seemed to be having a BPM problem. This was resolved and Operations resumed sending them normal beam at 11:03 AM. There was also a LCW problem, which required an access.

Operations established beam to the Meson MTest experiment at 11:05 AM.

At 11:56 AM, Operations shut off NuMI’s beam while experts further diagnosed their BPM problem. Operations resumed sending beam to NuMI at 12:34 PM so that experts could investigate with beam.

Pbar experts ended their reverse proton study at 2:46 PM.

At 3:22 PM, Operations reported MI beam problems and no beam to Switchyard (SY). Around 3 PM a MI expert restored the corrector settings from this morning to see if these changes were affecting NuMI. NuMI continued to run at low intensity.

At 3:56 PM, Operations reported that the interlocks dropped for the Meson M02 enclosure. Interlock technicians investigated.

Operations reported at 4:10 PM that CDF was turning their silicon detector back on (they had turned it off at some point during the morning). TeV experts had found a collimator that was in a bad position.

At 6:25 PM, Operations reported that there was still no SY beam. Operations rebooted the QXR at 7:23 PM. Once it “learned in,” they established beam.

The Recycler began preparing to stash at 7:30 PM.

Operations reported at 7:53 PM that NuMI was back to normal intensity. MI experts had found a bad clock cable, which resolved the SY problems and part of the NuMI problems. NuMI experts had found a problem with their orbit verifier program.

The stash ended. Pbar, MTest, and MCenter resumed normal operations at 8:29 PM.

At 9:17 PM, TRF1 tripped off.

Wednesday November 9

The midnight shift began with Operations monitoring store 4492 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

TeV experts began an end-of-store study at 4:35 AM.

At 4:36 AM, the NuMI Horn tripped off. Operator had it reset and back on ten minutes later.

At 5:08 AM, the NuMI target helium flow alarmed. Operators thought this was due to the target being off.

MTest experimenters reported at 5:08 AM that they were done with their beam.

Operations terminated store 4492 at 6:13 AM.

Operations prepared the CDF and D0 collision halls for access at 6:21 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster, Pbar, and MI to conduct studies.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/9/2005   email Fermilab